OpenStudy (anonymous):

solutions? (with work) 2=6p-8-5p -6y + 14 + 4y=32 5d-d-2d+8-3d=0 6/7 x -8=7

OpenStudy (anonymous):

2= 6p-8-5p 2+8=6p-5p 10=p p=10

OpenStudy (anonymous):

-6y+14+4y=32 -6y+4y=32-14 -2y=18 2y=-18 y=-18/2 y=-9

OpenStudy (anonymous):

5d-d-2d+8-3d=0 5d-d-2d-3d=-8 5d-6d=-8 -d=-8 d=8

OpenStudy (anonymous):

6/7 x -8=7 6/7 x =7+8 6/7 x =15 7/6(6/7 X)=7/6(15) X= 105/6 = 35/2= 17.5 I hope that will help

OpenStudy (anonymous):

Oh My Goodness, thank you so much !